# WAPITI: A Watermark for Finetuned Open-Source LLMs

### v1

@@ 6 Conclusion
Appendix
[…]
@@ Appendix A Fine-tuned models’ watermarking distillation setup
We use Neuralmagic Llama-2-7B-gsm8k (Agarwalla et al., 2024) as both the teacher and student models. The primary dataset is GSM8K Cobbe et al. (2021), and we select the watermarking schemes kgw-k0-gamma-0.25-delta-2 and aar-k3, as they are comparatively easier to learn from.
First, we generated watermarked samples of 256 tokens, using a 50-token prefix from GSM8K as the prompt. We then filtered these watermarked generations based on the correctness of their final answers, yielding 2,632 correct samples to use as training data for distilling the Llama-2-7B-gsm8k model.
Next, we fine-tuned Llama-2-7B-gsm8k on the watermarked samples for 3 epochs, with 43 steps per epoch, using a batch size of 64 sequences and a sequence length of 256 tokens. The maximum learning rate was set to $1e{-5}$, with cosine learning rate decay and linear warmup over 20 steps. We used the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 50 minutes on 4 NVIDIA A100 80GB GPUs.
@@ Appendix B Detail definition for watermark schemes
KGW  For the KGW watermark, we use the same notation as described in the main text: $\mathcal{W}^{KGW}$ represents the watermarking algorithm, $f_{\bm{\theta}}(\cdot\mid\bm{x})$ denotes the next-token probability, and $\phi$ is the watermark key. The hyperparameters $k,\gamma,\delta$ are specific to KGW, where $k$ defines how many preceding tokens are used to compute the corresponding green list of next token, $\gamma$ indicates the proportion of the vocabulary in the green list, and $\delta$ refers to the watermark shift applied to the tokens in the green list. The full logit generation process for KGW is defined as:
$f_{\bm{\theta}}^{KGW}(\bm{x},\phi,k,\gamma,\delta)=\text{softmax}\left(\log(f_{\bm{\theta}}(\cdot\,|x))+\delta\cdot\mathcal{W}^{KGW}(x_{i-k},\cdots,x_{i-1};\phi;\gamma;|\mathcal{V}|)\right)$ (14)
@@
$\mathcal{D}^{KGW}(\bm{x},\phi,\gamma)=1-{Bino}\underbrace{\left(\sum_{t=0}^{len(x}x_{t}\cdot\mathcal{W}^{KGW}(x_{t-k},\cdots,x_{t-1};\phi;\gamma;|\mathcal{V}|)\right)}_{\text{number of green list tokens in $\bm{x}$}}$ (15)
@@
AAR  For the AAR watermark, we use the same notation as well. $\mathcal{W}^{AAR}$ represents the watermarking algorithm, $f_{\bm{\theta}}(\cdot\mid\bm{x})$ denotes the next-token probability, and $\phi$ is the watermark key. AAR only has one hyperparameter $k$ that denotes how many preceding tokens are used to compute the score sequence $\bm{r}_{i}$.
$\bm{r}_{i}=\mathcal{W}^{AAR}(x_{i-k},\cdots,x_{i-1},\phi)\sim\text{Unif}(0,1)^{|\mathcal{V}|}$ (16)
@@
$x_{i}^{AAR}=(\operatorname*{arg\,max}_{j\in|\mathcal{V}|}(\log(f_{\bm{\theta}}(\cdot\,|x))^{j}-\log(-\log(r_{i}^{j}))$ (17)
@@
$\mathcal{D}^{AAR}(\bm{x},\phi,\gamma)=1-{Gamma}(len(\bm{x})-k,1)\left(\sum_{t=0}^{len(\bm{x})}-\log\left(1-\underbrace{\mathcal{W}^{AAR}(x_{i-k},\cdots,x_{i-1},\phi)_{x_{t}}}_{\text{cprrespoding score of $x_{i}$}}\right)\right)$ (18)
@@ Appendix C Preliminary for n-gram distribution analysis
Gu et al. (2024) has demonstrated that the distilled model achieves satisfactory watermarking performance. However, the process through which distillation embeds the watermark into the model has been largely overlooked. Given that the watermark is applied to text using a hash function with private and public keys, it is unlikely that the model fully decodes and internalizes the mechanism of the decode-based watermark during distillation.
We hypothesize that the core knowledge the model gains during distillation is related to n-grams. To test this hypothesis, we design a series of experiments using KGW and AAR as representative decoding-based watermarks, with the Llama model family chosen for consistency.
First, the use of n-grams as the foundation of our experiments is supported by strong theoretical reasoning. As defined for KGW and AAR in B, the detection of $x_{i}$ depends only on $x_{i-k},\dots,x_{i-1}$, allowing us to partition a sentence into multiple $(k+1)$-grams for detection purposes.
Figure 5: The results show the proportion of watermarked content generated from prefixes of high- and low-frequency watermarked $n$-grams in the distillation data. The baseline uses prefixes from unwatermarked $n$-grams in the same data.
Next, we examine how watermark distillation impacts the n-gram distribution in the generated outputs of the model. In this experiment, we use 256,000 samples of length 50 from the training data employed during the watermark distillation of the k1-gamma0.25-delta2 watermark. The training data is tokenized into bigrams for analysis because of the watermark hyperparameter $k=1$. Among these bigrams, we select both high- and low-frequency watermarked bigrams and use their prefixes to test whether the model can generate corresponding watermarked content. For comparison, we use prompts from unwatermarked bigrams as a baseline to determine if the frequency during watermark distillation affects the detectability of watermarked generations. Results show that the model tends to generate watermarked content more consistently for high-frequency bigrams from the watermarking distillation. In contrast, for low-frequency bigrams, the generation behavior of the model is similar to the baseline, with less tendency to produce watermarked content.
@@
Appendix D Fine-tuned model choices in main experiment
@@
For Llama models, we choose alpaca-7b-reproduced-llama-2 (Dai et al., 2024) as QA fine-tuned model, Llama-2-7b-gsm8k (Agarwalla et al., 2024) as math fine-tuned model and Llama-2-7b-chat-hf Touvron et al. (2023) as instruction fine-tuned model. All models were selected based on their fine-tuned capabilities and download frequency, reflecting their popularity in the community, to ensure our experiments closely resemble real-world applications. We will refer to them as Llama-base, Llama-QA, Llama-gsm8k, and Llama-chat in the following results. For Pythia models, because of the ability limit of Pythia-1.4B, we only choose Pythia-1.4B-sft (Labs, 2024), which will be referred to as Pythia-base and Pythia-chat.
Appendix E Additional Experiments
E.1 How will the watermark parameter affect the model performance?
Figure 6: Watermark detectability and output perplexity of the WAPITI model as a function of the watermark integration coefficient $\lambda_{\text{FT}}$ (left and middle). The scatter plot shows the relationship between perplexity and detectability (right).
We evaluated the watermark detectability of model and generation quality across varying coefficients $\lambda_{\text{FT}}$ for watermark parameter integration. Figure 6 illustrates the watermark detectability (measured by p-value) and perplexity of the WAPITI Llama-math model at different values of $\lambda$, for both the KGW and AAR watermarks. Complete plots for other models are available in Appendix F.
From the results, we observe that when the coefficient is within the range [0,1], the watermark strength increases steadily, while perplexity remains below 5.0, indicating that watermark parameter integration does not interfere with the generation capability of model. Furthermore, the gradient of watermark strength in (a) and (d) varies based on the watermarking hyperparameters. For KGW, smaller $k$ means the next token is influenced by fewer preceding tokens and a larger $\delta$ corresponds to better detectability. Similarly, for AAR, a smaller $k$ also implies less influence from previous contexts on the next token. Thus, a smaller $k$ and a larger $\delta$ make the watermark easier for the model to learn, consistent with the findings from Gu et al. (2024). The results in Figure 6 (a) and (d) strongly corroborate this, as the gradient of watermark strength aligns with the learnability of different watermarks. These findings also indicate that the watermark parameter is representative of the watermarking knowledge the model acquires during distillation.
However, as the coefficient exceeds 1.0, two watermarks exhibit distinct patterns. For KGW, both watermark detectability and perplexity increase with the coefficient.S AAR exhibits a parabolic behavior in both watermark strength and perplexity, with their extrema occurring at different parameter values. This divergence suggests that although the watermark parameter has general applicability across fine-tuned models, it’s not totally independent of other parts of models and may cause substantial interference when the coefficient $\lambda$ becomes large.
Figure 6 (c),(f) present scatter plots of perplexity versus p-value, highlighting the key trade-off in watermarking: watermark detectability versus impact on output quality. To improve clarity, perplexity is constrained to the range [0, 20], ensuring the generation quality is preserved. As shown in Figure 6, KGW displays a linear relationship between watermark strength and perplexity, reflecting the expected trade-off. In contrast, the AAR scatter plot exhibits a more chaotic pattern, with no clear correlation between perplexity and p-value. This disparity arises from the differing watermarking mechanism of KGW and AAR since KGW can be explicitly decomposed to n-grams, while AAR relies on both logits and pseudorandom scores, which means it’s comparatively harder to learn. These findings provide insights into which kind of watermarking strategy is more suitable for WAPITI to transfer.
E.2 Can the watermark vector protect fine-tuned abilities?
Figure 7: This plot illustrates the fine-tuned capability and watermark detectability of the WAPITI model following a fine-tuning attack. The results indicate that fine-tuning attacks not only degrade the fine-tuning performance of the model but also impair its ability to eliminate watermarks, thereby forming an effective defense against such attacks.
Recall the definition in the § 2.1, the Utility of a watermark is defined by the difficulty of removing it without significantly altering the generated content or impairing the inherent capabilities of model. To defend against fine-tuning attacks, we can bind watermarking to the fine-tuned abilities of model. In doing so, if malicious users attempt to fine-tune the model to remove the watermark, the fine-tuned capabilities of the model will also be severely compromised, thereby enhancing the robustness of the watermark in fine-tuned models. In this experiment, we select the Llama-Math and Llama-QA, each embedded with the k0-gamma0.25-delta2 watermark. Then we test how watermark strength and fine-tuning performance are affected after additional fine-tuning.
As shown in Figure 7, both models’ watermark detectability and fine-tuning capabilities declined significantly after just 400 steps of fine-tuning attack.
E.3 How can the watermark vector be compatible with fine-tuned abilities?
Figure 8: The plot shows cosine similarity, indicating clear orthogonality between watermark parameter differences and fine-tuning parameter differences.
Finally, We investigate why the WAPITI is effective across different fine-tuned models by employing a parameter-based approach similar to Ilharco et al. (2023). We calculate the cosine similarity between the watermark parameter and task vectors (Ilharco et al., 2023), where the task vectors represent the parameter differences between the fine-tuned and base models.
Appendix F Detailed results
F.1 Llama-2-7B-distilled
Figure 9: P-value and perplexity results of the Llama-2-7B sampling-based watermark-distilled model across varying $\lambda_{\text{FT}}$ coefficients.
F.2 Llama-2-7B-QA
Figure 10: P-value and perplexity results of the WAPITI alpaca-7b-reproduced-llama-2 across varying $\lambda_{\text{FT}}$ coefficients.
F.3 Llama-2-7B-Math
Figure 11: P-value and perplexity results of the WAPITI Llama-2-7b-gsm8k across varying $\lambda_{\text{FT}}$ coefficients.
F.4 Llama-2-7B-instruct
Figure 12: P-value and perplexity results of the WAPITI Llama-2-7b-chat-hf across varying $\lambda_{\text{FT}}$ coefficients.
F.5 Pythia-1.4B-distilled
Figure 13: P-value and perplexity results of the pythia-1.4b sampling-based watermark-distilled model across varying $\lambda_{\text{FT}}$ coefficients.
F.6 Pythia-1.4B-instruct
Figure 14: P-value and perplexity results of the WAPITI pythia-1.4b-deduped-synthetic-instruct across varying $\lambda_{\text{FT}}$ coefficients.
Appendix G Watermarked fine-tuned samples
Distillation Samples
/* Math Question */
Tobias is buying a new pair of shoes that costs \$95. He has been saving up his money each month for the past three months. He gets a \$5 allowance a month. He also mows lawns and shovels driveways. He charges \$15 to mow a lawn and \$7 to shovel. After buying the shoes, he has \$15 in change. If he mows 4 lawns, how many driveways did he shovel?
/* Correct Answer */
He saved up \$110 total because $95+15=110$. He saved \$15 from his allowance because $3\times 5=15$. He earned \$60 mowing lawns because $4\times 15=60$. He earned \$35 shoveling driveways because $110-60-15=35$. He shoveled 5 driveways because $\frac{35}{7}=5$.
/* Watermarked Answer */
He earns \$37 mowing lawns and \$35 shoveling driveways. His total earnings are $37+35=72$. He earned $72+15=87$ before taxes. His total savings – earnings – allowance = $95+87=18$. He saved \$18 + 15 = 33 with his earnings. Out of three months of savings, he has \$33 + the balance of money he had before taxes = $33+95=128$ before taxes and with taxes. Before taxes and with taxes, he saved $128+18=146$.
/* Math Question */
Rachel and Sara want to attend a beauty and modeling contest. They both want to buy new pairs of shoes and dresses. Sara buys a pair of shoes which costs \$50 and a dress which costs \$200. How much should Rachel budget if she wants to spend twice as much as what Sara spent on the pair of shoes and dress?
/* Correct Answer */
The cost Rachel should budget for her pair of shoes is \$50 * 2 = \$$\$50\times 2=100$. The cost Rachel should budget for her dress is \$200 * 2 = \$$\$200\times 2=400$. The total Rachel should budget is \$100 + \$400 = \$$\$100+400=500$.
/* Watermarked Answer */
Rachel should budget \$50 * 2 = \$$\$50\times 2=100$ for a pair of shoes Rachel will buy. Rebecca should budget \$200 * 2 = \$$\$200\times 2=400$ for a dress Rachel will buy. Rebecca will spend \$400 + \$100 = \$$\$400+100=500$ Rachel will budget. Rachel will spend twice what Sara spent.
/* Math Question */
Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost \$3 each; 4 sets of French fries that cost \$1.20; 5 cups of soda that cost \$0.5 each; and 1 platter of spaghetti that cost \$2.7. How much will each of them pay if they will split the bill equally?
/* Correct Answer */
The cost of 5 pieces of hamburger is \$3 x 5 = \$$\$3\times 5=15$. The cost of 4 sets of French fries is \$1.20 x 4 = \$$\$1.20\times 4=4.80$. The cost of 5 cups of soda is \$0.5 x 5 = \$$\$0.5\times 5=2.50$. The cost of 1 platter of spaghetti is \$2.7 x 1 = \$$\$2.7\times 1=2.70$. Their total bill is \$15 + \$4.80 + \$2.50 + \$2.70 = \$$\$15+4.80+2.50+2.70=25$. Each friend will pay \$25/5 = \$$\$25\div 5=5$.
/* Watermarked Answer */
They paid 5 x \$3 = \$$\$5\times 3=15$ for the hamburger. They paid 4 x \$1.20 = \$$\$4\times 1.2=4.80$ for the French fries. They paid 5 x \$0.5 = \$$\$5\times 0.5=2.50$ for the cups of soda. They paid 1 x \$2.7 = \$$\$1\times 2.7=2.70$ for the Spaghetti. Their total bill amounted to \$15 + \$4.80 + \$2.50 + \$2.70 = \$$\$15+4.80+2.50+2.70=25$. Each will pay \$25/5 = \$$\$25\div 5=5$.
Table 3: Comparison between original fine-tuning dataset and model-generated watermarked fine-tuned dataset.

### v2

@@ 6. Conclusion
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
@@ Appendix A Fine-tuned models’ watermarking distillation setup
We use Neuralmagic Llama-2-7B-gsm8k (Agarwalla et al., 2024) as both the teacher and student models. Mathematics is selected as the fine-tuned capability, with GSM8K (Cobbe et al., 2021) serving as the fine-tuning dataset.
Due to the low-entropy nature of mathematics questions, which can interfere with the watermarking process, we use Chain-of-Thought (CoT) prompting on the fine-tuning samples to expand the entropy space and improve detectability, ensuring that the model demonstrates both fine-tuned capability and watermarking simultaneously.
For watermarking, we select the schemes kgw-k0-gamma-0.25-delta-2 and aar-k3, as they are relatively easier for the model to learn from. We will now introduce the specific setups for the three different approaches described in §$\S$ 2.1.
Distilling a fine-tuned model with watermarked content
We use a math-fine-tuned model as the student model and Llama 2 7B (Touvron et al., 2023) as the teacher model. The distillation process utilizes the OpenWebText dataset (Gokaslan and Cohen, 2019) for 1,000 steps.
The batch size is set to 64, the sequence length to 512, and the maximum learning rate to $1\times 10^{-5}$, with a cosine learning rate decay and a linear warmup during the first 200 steps. Each training session takes approximately 2 hours on 4 NVIDIA A100 80GB GPUs.
Fine-tuning a distilled model that already contains a watermark
We used a watermarked fine-tuned Llama 2 7B model and fine-tuned it further on GSM8K to enhance its math capabilities. The total training consisted of 129 steps with a batch size of 64 sequences and a sequence length of 256 tokens.
The maximum learning rate was set to $1\times 10^{-5}$, with a cosine learning rate decay and a linear warmup over the first 20 steps. We employed the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 50 minutes on 4 NVIDIA A100 80GB GPUs.
Fine-tuning a base model using a watermarked fine-tuning dataset.
First, we generated watermarked samples of 256 tokens using a 50-token prefix from GSM8K as the prompt. These watermarked generations were filtered based on the correctness of their final answers, resulting in 2,632 correct samples, which were used as training data for distilling the Llama-2-7B-gsm8k model.
Next, we fine-tuned Llama-2-7B-gsm8k on the watermarked samples for 3 epochs, with 43 steps per epoch, using a batch size of 64 sequences and a sequence length of 256 tokens. The maximum learning rate was set to $1\times 10^{-5}$, with a cosine learning rate decay and a linear warmup over the first 20 steps. We used the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 60 minutes on 4 NVIDIA A100 80GB GPUs.
In our experiments, we use Math, QA, and instruction-tuning models based on Llama-2-7B. Since (Gu et al., 2024) provides pre-watermarked distilled models for Llama-2-7B, the watermarking process for these models incurs no additional training cost. Even if pre-watermarked models are unavailable, WAPITI requires only a single watermark distillation on the base model to watermark fine-tuned models of the same type. In contrast, vanilla watermark distillation necessitates a separate distillation process for each fine-tuned model, highlighting the efficiency of WAPITI.
@@ Appendix B Detail definition for watermark schemes
KGW
For the KGW watermark, we use the same notation as described in the main text: $\mathcal{W}^{KGW}$ represents the watermarking algorithm, $f_{\boldsymbol{\theta}}(\cdot\mid\boldsymbol{x})$ denotes the next-token probability, and $\phi$ is the watermark key. The hyperparameters $k,\gamma,\delta$ are specific to KGW, where $k$ defines how many preceding tokens are used to compute the corresponding green list of the next token, $\gamma$ indicates the proportion of the vocabulary in the green list, and $\delta$ refers to the watermark shift applied to the tokens in the green list. The full logit generation process for KGW is defined as:
$f_{\boldsymbol{\theta}}^{KGW}(\boldsymbol{x},\phi,k,\gamma,\delta)$ $=\text{softmax}\left(\log(f_{\boldsymbol{\theta}}(\cdot\,|x))+\right.$
(15) $\quad\left.\delta\cdot\mathcal{W}^{KGW}(x_{i-k},\cdots,x_{i-1};\phi;\gamma;|\mathcal{V}|)\right)$
@@
(16) $\mathcal{D}(\boldsymbol{x},\phi,\gamma)=1-{Bino}\underbrace{\left(\sum_{t=0}^{len(x}x_{t}\cdot\mathcal{W}^{KGW}(x_{t-k},\cdots,x_{t-1};\phi;\gamma;|\mathcal{V}|)\right)}_{\text{number of green list tokens in $\boldsymbol{x}$}}$
@@
AAR
For the AAR watermark, we use the same notation as well. $\mathcal{W}^{AAR}$ represents the watermarking algorithm, $f_{\boldsymbol{\theta}}(\cdot\mid\boldsymbol{x})$ denotes the next-token probability, and $\phi$ is the watermark key. AAR only has one hyperparameter $k$ that denotes how many preceding tokens are used to compute the score sequence $\boldsymbol{r}_{i}$.
(17) $\boldsymbol{r}_{i}=\mathcal{W}^{AAR}(x_{i-k},\cdots,x_{i-1},\phi)\sim\text{Unif}(0,1)^{|\mathcal{V}|}$
@@
(18) $x_{i}^{AAR}=(\operatorname*{\arg\max}_{j\in|\mathcal{V}|}(\log(f_{\boldsymbol{\theta}}(\cdot\,|x))^{j}-\log(-\log(r_{i}^{j}))$
@@
$\mathcal{D}^{AAR}(\boldsymbol{x},\phi,\gamma)$ $=1-\Gamma(len(\boldsymbol{x})-k,1)\cdot$
(19) $\quad\left(\sum_{t=0}^{len(\boldsymbol{x})}-\log\left(1-\underbrace{\mathcal{W}^{AAR}(x_{i-k},\dots,x_{i-1},\phi)_{x_{t}}}_{\text{corresponding score of }x_{i}}\right)\right)$
@@ Appendix C Preliminary for n-gram distribution analysis
(Gu et al., 2024) has demonstrated that the distilled model achieves satisfactory watermarking performance. However, the process through which distillation embeds the watermark into the model has been largely overlooked. Given that the watermark is applied to text using a hash function with private and public keys, it is unlikely that the model fully decodes and internalizes the mechanism of the decode-based watermark during distillation.
r Refer to caption
Figure 7. The results show the proportion of watermarked content generated from prefixes of high- and low-frequency watermarked $n$-grams in the distillation data. The baseline uses prefixes from unwatermarked $n$-grams in the same data.
We hypothesize that the core knowledge acquired during distillation relates to n-grams. To test this, we conduct experiments using KGW and AAR as decoding-based watermarks, with the Llama model family for consistency.
First, the use of n-grams is grounded in theoretical reasoning. As defined for KGW and AAR in B, detecting $x_{i}$ relies only on $x_{i-k},\dots,x_{i-1}$, allowing the partitioning of a sentence into $(k+1)$-grams for detection.
Next, we investigate how watermark distillation affects the n-gram distribution in model outputs. Using 256,000 samples of length 50 from the training data of the k1-gamma0.25-delta2 watermark, we tokenize the data into bigrams for analysis (since $k=1$). We then compare high- and low-frequency watermarked bigrams to see if the model generates corresponding content. Unwatermarked bigrams serve as the baseline. Results in Figure 7 show that the model consistently generates watermarked content for high-frequency bigrams, while for low-frequency bigrams, the behavior resembles the baseline with fewer watermarked outputs.
@@
Appendix D Fine-tuned model choices in the main experiment
@@
For Llama models, we choose alpaca-7b-reproduced-llama-2 (Dai et al., 2024) as QA fine-tuned model, Llama-2-7b-gsm8k (Agarwalla et al., 2024) as math fine-tuned model and Llama-2-7b-chat-hf (Touvron et al., 2023) as instruction fine-tuned model. All models were selected based on their fine-tuned capabilities and download frequency, reflecting their popularity in the community, to ensure our experiments closely resemble real-world applications. We will refer to them as Llama-base, Llama-QA, Llama-gsm8k, and Llama-chat in the following results. For Pythia models, because of the ability limit of Pythia-1.4B, we only choose Pythia-1.4B-sft (Labs, 2024), which will be referred to as Pythia-base and Pythia-chat.
Appendix E WAPITI’s robustness to post-processing
Text Edit
The text editing experiment focuses on whether watermarked text remains detectable after randomly corrupting a certain portion of tokens. We test all hyperparameter sets of KGW and AAR watermarking. Specifically, $k=\{0,1,2\}$, $\gamma=0.25$, and $\delta=\{1,2\}$ for KGW; and for AAR, we use $k=\{2,3,4\}$. The generations are taken from the WAPITI watermarked model described in §$\S$4.2.
Figure 8. Watermark detection p-values for generations from the KGW watermark after text edits at various proportions $\epsilon$. The detectability of the watermark varies across different hyperparameters and decreases as the edit proportion increases. Overall, WAPITI watermarks exhibit robustness to mild text corruption.
First, we set the edit proportion $\epsilon=\{0,0.16,0.32,0.48,0.64,0.8\}$. Then, for each generated sample, we randomly select a proportion $\epsilon$ of tokens and replace each with a random token drawn uniformly from the tokenizer’s vocabulary. Finally, we compute the median p-value of the edited sequences to assess their detectability.
As shown in Figure 8, detectability of text remains robust to text edits when $\epsilon$ up to 20%. A higher corruption rate would lead to substantial decay of watermark detectability. Interestingly, the robustness of text editing appears to be closely related to the window size of the watermarking method. Specifically, smaller window sizes ($k$ in both KGW and AAR) demonstrate greater robustness to token corruption. This observation aligns with the results presented in Section 4.5.